A person deposited Rs 50,000 in bank for 2 years at the rate of 10% p.a. compound interest compounded annually. But after one year, Bank has changed the policy and decided to pay semi annual compound interest at the same rate. What is the percentage difference between compound interests of first and second year?

Grade:10

1 Answers

Arun
4 years ago
Principle Amount , p= Rs.55000
Rate of interest per annum, R = 10%
 
First year interest = p×(R/100) = 55000×(10/100) = Rs. 5500
First year Amount = Rs.55000+Rs.5500 = Rs.60500
 
Second year compounded amount = 60500×[ 1+ (10)/(2×100) ]2 = Rs.66701.25  
 
Hence second year compound interest = 66701.25 - Rs.60500 = Rs.6201.25 
 
percentage of second year compound interest = (6201.25/60500)×100 = 10.25 %
 
Hence difference in percentage of compound interest = 10.25 - 10 = 0.25%